Using table II to obtain the areas under the standard normal curve required in exercises 6.59-6.66. Sketch a standard normal curve and shade the area of interest in each problem.

6.64 determine the area under the standard normal curve that lies between a. -0.88 and 2.24 b-2.5 and -2 c. 1.48 and 2.72 d -5.1 and 1
Answer #1

a)

P(-0.88 <Z<2.24)=0.9875- 0.1894 =0.7981

b)

P(-2.5 <Z<-2 )=0.0228-0.0062 =0.0166

c)

P(1.48 <Z< 2.72)=0.9967-0.9306 =0.0661

d)

P(-5.1 <Z <1)=0.8413-0.000=0.8413